Kathryn is 8 years older than her sister Rosemary. Eight years ago Kathryn was 3 times as old as Rosemary. How old is each now?

Answer by macston(5194)

.
R=Rosemary's age; K=Kathryn's age=R+8yrs
.
K-8yrs=3(R-8yrs) Substitute for K
(R+8yrs)-8yrs=3(R-8yrs)
R=3R-24yrs
-2R=-24yrs
R=12 yrs ANSWER 1: Rosemary is 12 years old.
K=R+8yrs=12yrs+8yrs=20 yrs ANSWER 2: Kathryn is 20 years old.
.
CHECK:
K-8yrs=3(R-8yrs)
20yrs-8yrs=3(12yrs-8yrs)
12yrs=3(4yrs)
12yrs=12yrs